# Image Captioning Evaluation in the Age of Multimodal LLMs: Challenges and Future Perspectives

### v1

@@ Abstract
The evaluation of machine-generated image captions is a complex and evolving challenge. With the advent of Multimodal Large Language Models (MLLMs), image captioning has become a core task, increasing the need for robust and reliable evaluation metrics. This survey provides a comprehensive overview of advancements in image captioning evaluation, analyzing the evolution, strengths, and limitations of existing metrics. We assess these metrics across multiple dimensions, including correlation with human judgment, ranking accuracy, and sensitivity to hallucinations. Additionally, we explore the challenges posed by the longer and more detailed captions generated by MLLMs and examine the adaptability of current metrics to these stylistic variations. Our analysis highlights some limitations of standard evaluation approaches and suggest promising directions for future research in image captioning assessment.
@@ 1 Introduction
Training strategies have evolved, incorporating supervised learning integrating both cross-entropy loss and reinforcement learning using self-critical sequence training Rennie et al. (2017). Some datasets have supported the development of captioning models, including COCO Lin et al. (2014), Conceptual Captions Sharma et al. (2018), and nocaps Agrawal et al. (2019), providing varying diversity and complexity.
@@
Recent advances have shifted towards learnable metrics leveraging vision-and-language pre-trained models. CLIP-S Hessel et al. (2021) introduce a contrastive approach to measure image-caption alignment based on CLIP embeddings. Fine-grained metrics like BRIDGE Sarto et al. (2024a) and HiFi-S Yao et al. (2024) capture localized semantics and hierarchical structures. LLM-based metrics such as CLAIR Chan et al. (2023) and FLEUR Lee et al. (2024) integrate reasoning and explanation into caption evaluation.
With the rise of MLLMs as image description generators, the nature of captions has changed, as these models generate captions that differ in length, specificity, and abstraction. Conventional evaluation metrics may no longer reliably assess caption quality in this new paradigm, posing new challenges to the evaluation of generated captions. Following this insight, this survey provides a comprehensive analysis of existing captioning evaluation metrics, from traditional rule-based approaches to modern learnable methods and LLM-driven techniques. We explore their evolution, strengths, and limitations while emphasizing the growing need for robust evaluation frameworks that can adapt to the stylistic diversity introduced by MLLMs. Additionally, we present an in-depth experimental study assessing metrics across multiple dimensions, including correlation with human judgment, ranking accuracy, and sensitivity to hallucinations. Our analysis highlights the performance gaps of conventional metrics and showcase promising directions for future research.
@@ 2.2 Learnable Metrics
TIGEr Jiang et al. (2019), on the other hand, focuses on the integration of the image as one of the inputs for computing an evaluation score. Building upon a text-image grounding model Lee et al. (2018), it allows evaluating caption quality not only based on how well a caption represents image content but also on how well machine-generated captions match human-generated ones.
Following this line and inspired by BERT-S, ViLBERT-S Lee et al. (2020) computes cosine similarity between token embeddings for reference and candidate sentences. However, different from BERT-S, the token embedding is computed with the consideration of image contexts. On a different line, FAIEr Wang et al. (2021) employs the scene graph as a bridge between the image and textual modality. Specifically, it fuses scene graphs of the image and references as a union scene graph and compares it with the scene graph of generated captions.
@@
Available in both reference-free and reference-based versions, HICE-S Zeng et al. (2024) highlights the limitations of CLIP-based metrics, which primarily assess global image-text compatibility but often struggle with detecting local textual hallucinations and maintaining sensitivity to small visual objects. To overcome these challenges, HICE-S introduces a hierarchical scoring mechanism that leverages the SAM model Kirillov et al. (2023) to generate masks for localized visual regions and corresponding textual phrases. These are then processed through a modified version of the CLIP architecture Sun et al. (2024), enhancing the model ability to capture fine-grained visual-semantic relationships.
@@ 3.1 Correlation with Human Judgment
We analyze the correlation of metrics with human judgment, highlighting their varying behaviors. Experimental results are reported on standard captioning evaluation datasets, including Flickr8k-Expert, Flickr8k-CF, and Composite, along with the more recent Polaris and Nebula datasets.
Datasets. Flickr8k-Expert Hodosh et al. (2013) contains 17k annotations for 5,664 images, with each pair scored from 1 (no correlation) to 4 (accurate depiction). Flickr8k-CF Hodosh et al. (2013) provides 145k binary quality judgments for 48k pairs across 1,000 unique images, using the mean proportion of “yes” votes as the alignment score. The Composite dataset Aditya et al. (2015) includes 12k human ratings for image-caption pairs (with around 4k unique images), assessed on a 1–5 scale. However, these datasets lack model-generated captions, leading to a domain gap when applying them to train evaluation metrics. To address this limitation, the Polaris dataset Wada et al. (2024) introduces 131k human judgments from 550 evaluators (i.e., around ten times larger than previous datasets) covering both human written and machine-generated captions from ten image captioning models. Expanding on Polaris, the Nebula dataset Matsuda et al. (2024) includes 32,978 images with human judgments from 805 annotators. In line with prior studies Hessel et al. (2021); Wada et al. (2024), we use Kendall’s $\tau_{b}$ for Flickr8k-CF and Kendall’s $\tau_{c}$ for the other datasets.
@@ 3.4 System-level Correlation
When evaluating captions generated by MLLMs interesting patterns emerge. If the captions remain similar in length to ones contained in COCO, rule-based metrics still perform effectively: CIDEr, for instance, assigns a score of 142.2 to captions generated by InstructBLIP. However, as caption length increases, these metrics exhibit a sharp decline in scores. Notably, when evaluating LLaVA-1.5, which produces significantly longer captions than COCO ones, CIDEr score drops dramatically to just 0.3. Qualitative results showing this trend are reported in Fig. 2. This decline is primarily due to the reliance of rule-based metrics on COCO-style reference captions. When comparing longer and more detailed MLLM-generated captions against these shorter and simpler references, these metrics often struggle to compute realistic scores, leading to misleading evaluations.
@@ 4 Future Directions and Open Problems
Benchmark Evolution. Traditional evaluation datasets mainly feature short captions similar to those in COCO, creating a gap with the longer, more detailed outputs of modern MLLMs. While recent efforts, like Polaris and Nebula, incorporate captions from standard models, they often maintain the concise style of COCO descriptions. A valuable direction would involve creating new benchmarks specifically designed to assess the quality of metrics on longer, richer captions that better reflect the MLLM outputs, addressing challenges like synonyms, paraphrases, and domain-specific terms.

### v2

@@ Abstract
The evaluation of machine-generated image captions is a complex and evolving challenge. With the advent of Multimodal Large Language Models (MLLMs), image captioning has become a core task, increasing the need for robust and reliable evaluation metrics. This survey provides a comprehensive overview of advancements in image captioning evaluation, analyzing the evolution, strengths, and limitations of existing metrics. We assess these metrics across multiple dimensions, including correlation with human judgment, ranking accuracy, and sensitivity to hallucinations. Additionally, we explore the challenges posed by the longer and more detailed captions generated by MLLMs and examine the adaptability of current metrics to these stylistic variations. Our analysis highlights some limitations of standard evaluation approaches and suggests promising directions for future research in image captioning assessment.
@@ 1 Introduction
Training strategies have also evolved, incorporating supervised learning integrating both cross-entropy loss and reinforcement learning using self-critical sequence training Rennie et al. (2017). Several benchmark datasets have facilitated the development of image captioning models, including COCO Lin et al. (2014), Conceptual Captions Sharma et al. (2018), and nocaps Agrawal et al. (2019), each providing varying levels of diversity and complexity.
@@
Recent advances have shifted towards learnable metrics leveraging vision-and-language pre-trained models. CLIP-S Hessel et al. (2021) introduces a contrastive approach to measure image-caption alignment based on CLIP embeddings. Fine-grained metrics like BRIDGE Sarto et al. (2024a) and HiFi-S Yao et al. (2024) capture localized semantics and hierarchical structures. LLM-based metrics such as CLAIR Chan et al. (2023) and FLEUR Lee et al. (2024) integrate reasoning and explanation into caption evaluation.
With the rise of MLLMs as image description generators, the nature of captions has changed, as these models generate captions that differ in length, specificity, and abstraction. Conventional evaluation metrics may no longer reliably assess caption quality in this new paradigm, posing new challenges to the evaluation of generated captions. Following this insight, this survey provides a comprehensive analysis of existing captioning evaluation metrics, from traditional rule-based approaches to modern learnable methods and LLM-driven techniques. We explore their evolution, strengths, and limitations while emphasizing the growing need for robust evaluation frameworks that can adapt to the stylistic diversity introduced by MLLMs. Additionally, we present an in-depth experimental study assessing metrics across multiple dimensions, including correlation with human judgment, ranking accuracy, and sensitivity to hallucinations. Our analysis highlights the performance gaps of conventional metrics and showcases promising directions for future research.
@@ 2.2 Learnable Metrics
TIGEr Jiang et al. (2019), on the other hand, focuses on the integration of the image as one of the inputs for computing an evaluation score. Building upon a text-image grounding model Lee et al. (2018), it allows evaluating caption quality not only based on how well a caption represents image content but also on how well machine-generated captions match human-generated ones. Following this line and inspired by BERT-S, ViLBERT-S Lee et al. (2020) computes cosine similarity between token embeddings for reference and candidate sentences. However, different from BERT-S, the token embedding is computed with the consideration of image contexts. On a different line, FAIEr Wang et al. (2021) employs the scene graph as a bridge between the image and textual modality. Specifically, it fuses scene graphs of the image and references as a union scene graph and compares it with the scene graph of generated captions.
@@
Available in both reference-free and reference-based versions, HICE-S Zeng et al. (2024) highlights the limitations of CLIP-based metrics, which primarily assess global image-text compatibility but often struggle with detecting local textual hallucinations and maintaining sensitivity to small visual objects. To overcome these challenges, HICE-S introduces a hierarchical scoring mechanism that leverages the SAM model Kirillov et al. (2023) to generate masks for localized visual regions and corresponding textual phrases. These are then processed through a modified version of the CLIP architecture Sun et al. (2024), enhancing the model’s ability to capture fine-grained visual-semantic relationships.
@@ 3.1 Correlation with Human Judgment
We analyze the correlation of metrics with human judgment, highlighting their varying behaviours. Experimental results are reported on standard captioning evaluation datasets, including Flickr8k-Expert, Flickr8k-CF, and Composite, along with the more recent Polaris and Nebula datasets.
Datasets. Flickr8k-Expert Hodosh et al. (2013) contains 17k annotations for 5,664 images, with each pair scored from 1 (no correlation) to 4 (accurate depiction). Flickr8k-CF Hodosh et al. (2013) provides 145k binary quality judgments for 48k pairs across 1,000 unique images, using the mean proportion of “yes” votes as the alignment score. The Composite dataset Aditya et al. (2015) includes 12k human ratings for image-caption pairs (with around 4k unique images), assessed on a 1–5 scale. However, these datasets lack model-generated captions, leading to a domain gap when applying them to train evaluation metrics. To address this limitation, the Polaris dataset Wada et al. (2024) introduces 131k human judgments from 550 evaluators (i.e., around ten times larger than previous datasets) covering both human-written and machine-generated captions from ten image captioning models. Expanding on Polaris, the Nebula dataset Matsuda et al. (2024) includes 32,978 images with human judgments from 805 annotators. In line with prior studies Hessel et al. (2021); Wada et al. (2024), we use Kendall’s correlation coefficient $\tau_{b}$ for Flickr8k-CF and Kendall’s $\tau_{c}$ for the other datasets.
@@ 3.4 System-level Correlation
When evaluating captions generated by MLLMs interesting patterns emerge. If the captions remain similar in length to the ones contained in COCO, rule-based metrics still perform effectively: CIDEr, for instance, assigns a score of 142.2 to captions generated by InstructBLIP. However, as caption length increases, these metrics exhibit a sharp decline in scores. Notably, when evaluating LLaVA-1.5, which produces significantly longer captions than COCO ones, CIDEr score drops dramatically to just 0.3. Qualitative results showing this trend are reported in Fig. 2. This decline is primarily due to the reliance of rule-based metrics on COCO-style reference captions. When comparing longer and more detailed MLLM-generated captions against these shorter and simpler references, these metrics often struggle to compute realistic scores, leading to misleading evaluations.
@@ 4 Future Directions and Open Problems
Benchmark Evolution. Traditional evaluation datasets mainly feature short captions similar to those in COCO, creating a gap with the longer, more detailed outputs of modern MLLMs. While recent efforts, like Polaris and Nebula, incorporate captions from standard models, they often maintain the concise style of COCO descriptions. A valuable direction would involve creating new benchmarks specifically designed to assess the quality of metrics on longer, richer captions that better reflect the MLLM outputs, addressing challenges like synonyms, paraphrases, and domain-specific terms.m